A mass of 5 kg is moving at 8.2 m/s in the positive x direction and collides in a perfectly elastic collision with a second mass which is not at rest. After the collision, the second mass is moving at 12.3 m/s in the +x direction. If the 5 kg mass stops after the collision, what is the amount of the second mass in kg?
Answer #1

we will conserve momentum,

5 ( 8.2) +m ( u) = m ( 12.3)

12.3 + u = 8.2 + 0 ( in a perfectly elastic collsion , velocity of approach is equal to velocity of recession)

u ( velocity of unknown mass) = 8.2- 12.3= -4.1 m/s

5(8.2) + m ( -4) = m ( 12.3)

41 = m ( 16.3)

m = 2.515 kg apprx